the length of a rectangle is six times its width. if the area of the rectangle is 384 yards squared find its perimeter
Answer by jankiz(13)

To make this problem easier, draw a rectangles. Label the length 6w because it is 6 times the width. Label width w. Then set up an equation that says length times width = area. This is the equation: 6w*w=384. Multiply, you get 6w^2=384, divide both sides by six, you get w^2=64, then take the square root of both sides, so sqrtw^2=sqrt64. x=8. Width is 8 and the length is 48! Then add all sides, 8+8+48+48=112.